1.     Suppose that the short-run production function of certain cut-flower firm is given by: Q= 4KL -0.6K2 -0.1L2 where Q is quantity of cut-flower produced, L is labour input and K is fixed capital input (K=5).

 a) Determine the average product of labour (APL) function.

b) At what level of labour does the total output of cut-flower reach the maximum?

c) What will be the maximum achievable amount of cut-flower production? 



1
Expert's answer
2021-02-03T20:22:50-0500

"Q= 4KL-0.6K^2-0.1L^2"

K= 5

Therefore by substituting K in the above function we get,

"Q= 4*5L-0.6(5)^2-0.1L^2"

"= 20L-15-0.1L^2"

By rearranging

"Q= -0. 1L^2+20L-15 .............. Equation-1"

Part (a) - Average product of labour function= "\frac{Q}{L}"

Therfore Equation "\frac{1}{L}" ,we get,

"APL= \frac{(-0.1L^2+ 20L-15)} {L}"

"APL = -0. 1L + 20-(\frac{15}{L})"

Part(b) - Now differentiate the Equation -1 with respect to labour(L)

"\frac{dQ}{dL}=\frac{ [d(-0.1L^2+20L-15) ]}{dL}"

= -0. 2L+20

Now according to first order condition,

"\frac{dQ}{dL}" =0

Therefore, -0. 2L+20= 0

L= 100

Part(c) - Now substitute the value of labour (L) in equation-1 to get maximum output,

We get,

"Q= -0. 1(100) ^2+20\times100-15"

Q = -1000+2000-15

Q= 985

Therefore maximum output will be 985 quantities.